Suppose the demand for crossing the Brooklyn Bridge is given by Q = 10,000 - 1000P. If the toll price is $3, what is the price elasticity of demand at this point? (Round up the nearest two decimal points. Don't include any negative signs in your answer.)
Answer #1

Answer

price elasticity of demand =(dQ/dP)*(P/Q)

dQ/dP=-1000 ............ the first differentiation of the demand curve concerning price by power rule.

Q=10000-1000*3=7000

Price elasticity of demand =(-1000)*(3/7000)

=-0.428571429

the price elasticity of demand is 0.43 (in absolute value)

the demand is inelastic.